# Linear orders on chainable continua

### v1

@@ 5. Order type of ultrafilter orders on Suslinean chainable continua
Let $X$ be a hereditarily decomposable chainable continuum. Then there exists a continuous and monotone surjection $f:X\to[0,1]$ such that if $g:X\to[0,1]$ is any other monotone and continous surjection, then there is a monotone continous surjection $m:[0,1]\to[0,1]$ such that $g=m\circ f$. Moreover, for every $t\in[0,1]$ we have $int(f^{-1}(t))=\emptyset$.
@@ 6.1. Order topology generated by a certain ultrafilter order on a Knaster continuum
Let $\mathcal{U}$ be a non-principal ultrafilter on $\mathbb{N}$. Let $\mathcal{D}=\{D_{n}\}_{n\in\mathbb{N}}$ (where $D_{n}=\{d_{i,n}\}_{i=1}^{k_{n}}$) be a sequence of chains covering the Knaster continuum such that for every $n$, $(0,0)\in d_{1,n}$ and $mesh(D_{n})\xrightarrow[]{n\to\infty}0$. It is easy to see that choosing such a sequence of chains is possible.
@@
Note that the composant of the Knaster continuum containing the point $(0,0)$ in the space $(K,\tau_{\mathcal{U}}^{\mathcal{D}})$ has the order type of the interval $[0,1)$ (and the point $(0,0)$ is the smallest point in the sense of the order $\leq_{\mathcal{U}}^{\mathcal{D}}$), and all the remaining composants in $(K,\tau_{\mathcal{U}}^{\mathcal{D}})$ have the order type of the open interval. This follows from the theorem 4.1 and from the fact that each component of a Knaster continuum is the sum of an ascending sequence of arcs.
Let us also observe that all components of a Knaster continuum are open in $(K,\tau_{\mathcal{U}}^{\mathcal{D}})$, which follows from the minimality of the point $(0,0)$, Lemma 4.2, and order types of the components.
@@
Let $\mathcal{U}$ be a nonprincipal ultrafilter on $\mathbb{N}$. Let $\mathcal{D}=\{D_{n}\}_{n\in\mathbb{N}}$ (where $D_{n}=\{d_{i,n}\}_{i=1}^{k_{n}}$) is such a sequence of chains covering the Knaster continuum, that for every $n$, $(0,0)\in d_{1,n}$ and $mesh(D_{n})\xrightarrow[]{n\to\infty}0$.
Then the Knaster continuum with the order topology $\tau_{\mathcal{U}}^{\mathcal{D}}$, generated by an ultrafilter order $\leq_{\mathcal{U}}^{\mathcal{D}}$, is homeomorphic to the disjoint sum of the topological spaces $X_{i}$:
@@
where $X_{0}$ is a space homeomomorphic to the interval $[0,1)$, corresponding to the arc component of the Knaster continuum containing the point $(0,0)$, and all other $X_{i}$ are homeomorphic to the open interval $(0,1)$ and correspond to the remaining arc components of the Knaster continuum.
@@ 7.1. Arc
The above observation implies that the identity function $id:(X,\tau)\to(X,\tau_{\mathcal{U}}^{\mathcal{D}})$ is continuous. From the fact that the identity is a continuous function defined on the compact space $X$ we conclude that $id$ is a homeomorphism. Since $(X,\tau_{\mathcal{U}}^{\mathcal{D}})$ is a linearly ordered space, then the space $(X,\tau)$, which is homeomorphic to it, is also linearly ordered. Using a well known fact that every separable linearly ordered continuum is homeomorphic to a closed interval [En, 6.3.2(b)], we obtain the thesis. ∎
@@ 7.2. Suslinean continua
By Theorem 5.2 there exists an order isomorphism $h:(X,\leq_{\mathcal{U}}^{\mathcal{D}})\to([0,1],\leq)$ which is a Borel isomorphism of the class (1,1) between $(X,\tau)$ and $(([0,1],\tau_{e})$. Then the map $H=h\times h:(X,\tau)^{2}\to(([0,1],\tau_{e})^{2}$ is of the first Baire class. The sets
@@ 7.3. The Knaster continuum
Let $\leq_{\mathcal{U}}^{\mathcal{D}}$ be any ultrafilter order on $K$. Let $C=K\cap l$, where $l$ is a line described as $x=\frac{1}{2}$. Let $l^{\prime}$ be a line desribed as $x=\frac{9}{20}$ and let $C^{\prime}=K\cap l^{\prime}$. Notice that $C$ and $C^{\prime}$ are homeomorphic to the Cantor set. For every $x\in C$ let $H_{x}$ be the unique semicircle from definition of $K$ (see Definition 6.1), s.t. $x\in H_{x}$ and let $x^{\prime}\in C^{\prime}$ be such a point in $H_{x}$, which lies on line $l^{\prime}$.
@@
Let $n\in\mathbb{N}$. For $x\in A_{n}$ let $I_{x}$ be the unique semicircle with the center at $(\frac{5}{2\cdot 3^{n+1}},0)$ contained in $K$, such that $I_{x}\cap H_{x}\neq\varnothing$. Let $J_{x}$ be the unique semicircle with the center at $(\frac{1}{2},0)$, contained in $K$ and such that $I_{x}\cap J_{x}\neq\varnothing$ and $I_{x}\cap H_{x}=\varnothing$.
@@
$\{s_{k_{i}}:0\leq i\leq m_{n}\}$ satysfying $k_{0}<k_{1}<k_{2}<...<n$ such that $s_{k_{0}}\circ...\circ s_{m_{n}}(B_{s})\subseteq A_{n}$. Apply $s_{n}\restriction_{A_{n}}$ to the set $s_{k_{0}}\circ...\circ s_{m_{n}}(B_{s})$ and notice that
@@
where $\pi_{1}$ is a projection onto the first coordinate. This projection restricted to the graph of $g$ is a homeomorphism, so from the fact that $M\cap Gr(g)$ is analytic (co-analytic) we obtain that $B=\pi_{1}(M\cap Gr(g))$ is also analytic (co-analytic), hence it has the proprty of Baire. A contradiction with Lemma 7.15. ∎
@@ 8. Questions
$x\leq_{\mathcal{U}}^{\mathcal{D}}y\iff h^{\prime}(x)\leq_{\mathcal{U}}^{(I_{i},f_{i})_{i=1}^{\infty}}h^{\prime}(y)$
@@
One can easily show that ultrafilter orders on the inverse limit homeomorphic to some chainable continuum (in the sense of Definition 3.5) are dense. We have a conjecture that ultrafilter orders defined using sequence of chains obtained from chainability of $X$ are also dense orders.
Question 8.2.
@@
Question 8.4.
Let $X$ be a chainable, indecomposable continuum. Is it true that every composant of $X$ is open in $(X,\tau_{\mathcal{U}}^{\mathcal{D}})$?
Question 8.5.
@@
By Corollary 6.5 we know that order topology generated by a certain ultrafilter order on the Knaster continuum is non-connected and non-compact. We would like to ask if order topology on any indecomposable chainable continuum also has those properties.
Question 8.6.
@@
Problem 8.7.

### v2

@@ 5. Order type of ultrafilter orders on Suslinean chainable continua
Let $X$ be a hereditarily decomposable chainable continuum. Then there exists a continuous and monotone surjection $f:X\to[0,1]$ such that if $g:X\to[0,1]$ is any other monotone and continuous surjection, then there is a monotone continuous surjection $m:[0,1]\to[0,1]$ such that $g=m\circ f$. Moreover, for every $t\in[0,1]$ we have $int(f^{-1}(t))=\emptyset$.
@@ 6. Ultrafilter orders on the Knaster continuum
Let us assume that $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ are distinct ultrafilters on $\mathbb{N}$.
@@
We will show that $\leq_{\mathcal{U}_{1}}^{(I_{i},f_{i})_{i=1}^{\infty}}\neq\leq_{\mathcal{U}_{2}}^{(I_{i},f_{i})_{i=1}^{\infty}}$ holds, i.e., ultrafilter orders on the inverse limit homeomorphic to the Knaster continuum, generated by ultrafilters $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$, are distinct.
@@ 6.1. Order topology generated by a certain ultrafilter order on a Knaster continuum
Let $\mathcal{U}$ be a non-principal ultrafilter on $\mathbb{N}$. Let $\mathcal{D}=\{D_{n}\}_{n\in\mathbb{N}}$ (where $D_{n}=\{d_{i,n}\}_{i=1}^{k_{n}}$) be a sequence of chains covering the Knaster continuum such that for every $n$, $(0,0)\in d_{1,n}$ and $mesh(D_{n})\xrightarrow[]{n\to\infty}0$. It is easy to see that selection of such a sequence of chains is possible.
@@
Note that the composant of the Knaster continuum containing the point $(0,0)$ in the space $(K,\tau_{\mathcal{U}}^{\mathcal{D}})$ has the order type of the interval $[0,1)$ (and the point $(0,0)$ is the smallest point in the sense of the order $\leq_{\mathcal{U}}^{\mathcal{D}}$), and all the remaining composants in $(K,\tau_{\mathcal{U}}^{\mathcal{D}})$ have the order type of the open interval. This follows from Theorem 4.1 and from the fact that each composant of a Knaster continuum is the sum of an ascending sequence of arcs.
Let us also observe that all composants of a Knaster continuum are open in $(K,\tau_{\mathcal{U}}^{\mathcal{D}})$, which follows from the minimality of the point $(0,0)$, Lemma 4.2, and order types of the composants.
@@
Let $\mathcal{U}$ be a non-principal ultrafilter on $\mathbb{N}$. Let $\mathcal{D}=\{D_{n}\}_{n\in\mathbb{N}}$ (where $D_{n}=\{d_{i,n}\}_{i=1}^{k_{n}}$) be such a sequence of chains covering the Knaster continuum, that for every $n$, $(0,0)\in d_{1,n}$ and $mesh(D_{n})\xrightarrow[]{n\to\infty}0$.
Then the Knaster continuum with the order topology $\tau_{\mathcal{U}}^{\mathcal{D}}$, generated by an ultrafilter order $\leq_{\mathcal{U}}^{\mathcal{D}}$, is homeomorphic to the disjoint sum of topological spaces $X_{i}$:
@@
where $X_{0}$ is a space homeomorphic to the interval $[0,1)$, corresponding to the arc component of the Knaster continuum containing the point $(0,0)$, and all other $X_{i}$ are homeomorphic to the open interval $(0,1)$ and correspond to the remaining arc components of the Knaster continuum.
@@ 7.1. Arc
The above observation implies that the identity function $id:(X,\tau)\to(X,\tau_{\mathcal{U}}^{\mathcal{D}})$ is continuous. From the fact that the identity is a continuous function defined on the compact space $X$ we conclude that $id$ is a homeomorphism. Since $(X,\tau_{\mathcal{U}}^{\mathcal{D}})$ is a linearly ordered space, the space $(X,\tau)$, which is homeomorphic to it, is also linearly ordered. Using a well known fact that every separable linearly ordered continuum is homeomorphic to a closed interval [En, 6.3.2(b)], we obtain the thesis. ∎
@@ 7.2. Suslinean continua
By Theorem 5.2 there is an order isomorphism $h:(X,\leq_{\mathcal{U}}^{\mathcal{D}})\to([0,1],\leq)$ which is a Borel isomorphism of the class (1,1) between $(X,\tau)$ and $([0,1],\tau_{e})$. Then the map $H=h\times h:(X,\tau)^{2}\to([0,1],\tau_{e})^{2}$ is of the first Baire class. The sets
@@
The following example shows that the converse of Theorem 7.4 does not hold.
Example 7.5.
Let $\mathcal{C}\subseteq[0,1]\subseteq\mathbb{R}$ be the standard Cantor set.
Let $X$ be a chainable continuum defined as:
$X=(\mathcal{C}\times[-1,1])\cup\Big\{\Big(t,\sin\Big(\frac{1}{dist(t,\mathcal{C})}\Big)\Big):t\in[0,1]\setminus\mathcal{C}\Big\}.$
Then $X$ is a non-Suslinean chainable continuum.
Notice that using a similar technique as in the proof of Theorem 4.10 we can obtain a sequence of chains $\mathcal{D}=\{D_{n}\}_{n\in\mathbb{N}}$ in $X$, such that $mesh(D_{n})\xrightarrow[]{n\to\infty}0$ and, for any non-principal ultrafilter $\mathcal{U}$, the ultrafilter order $\leq_{\mathcal{U}}^{\mathcal{D}}$ on $X$ is the same as lexicographic order $\leq_{lex}$ on $0,1]^{2}$, restricted to $X$ (i.e. $\leq_{\mathcal{U}}^{\mathcal{D}}~=~\leq_{lex}\restriction_{X}$). One can easily check that the lexicographic order $\leq_{lex}$ on $0,1]^{2}$ is both of type $F_{\sigma}$ and $G_{\delta}$ (in $0,1]^{4}$), hence $\leq_{lex}\restriction_{X}$ is also both of type $F_{\sigma}$ and $G_{\delta}$ (in $X^{2}$).
Notice that if $\tau_{\mathcal{U}}^{\mathcal{D}}$ is the order topology on the space $X$, generated by the ultrafilter order $\leq_{\mathcal{U}}^{\mathcal{D}}$ from Example 7.5, then the space $(X,\tau_{\mathcal{U}}^{\mathcal{D}})$ is compact (since each its subset has a supremum) and connected (since it is compact and has no gaps). Note also that the space $(X,\tau_{\mathcal{U}}^{\mathcal{D}})$ is non-metrizable, since it is compact and not ccc.
@@ 7.3. The Knaster continuum
Let $\leq_{\mathcal{U}}^{\mathcal{D}}$ be any ultrafilter order on $K$. Let $C=K\cap l$, where $l$ is a line described as $x=\frac{1}{2}$. Let $l^{\prime}$ be a line described as $x=\frac{9}{20}$ and let $C^{\prime}=K\cap l^{\prime}$. Notice that $C$ and $C^{\prime}$ are homeomorphic to the Cantor set. For every $x\in C$ let $H_{x}$ be the unique semicircle from definition of $K$ (see Definition 6.1), s.t. $x\in H_{x}$ and let $x^{\prime}\in C^{\prime}$ be such a point in $H_{x}$, which lies on line $l^{\prime}$.
@@
Let $n\in\mathbb{N}$. For $x\in A_{n}$ let $I_{x}$ be the unique semicircle with the center at $(\frac{5}{2\cdot 3^{n+1}},0)$ contained in $K$, such that $I_{x}\cap H_{x}\neq\varnothing$. Let $J_{x}$ be the unique semicircle with the center at $(\frac{1}{2},0)$, contained in $K$ and such that $I_{x}\cap J_{x}\neq\varnothing$ and $J_{x}\cap H_{x}=\varnothing$.
@@
$\{s_{k_{i}}:0\leq i\leq m_{n}\}$ satisfying $k_{0}<k_{1}<k_{2}<...<n$ such that $s_{k_{0}}\circ...\circ s_{m_{n}}(B_{s})\subseteq A_{n}$. Apply $s_{n}\restriction_{A_{n}}$ to the set $s_{k_{0}}\circ...\circ s_{m_{n}}(B_{s})$ and notice that
@@
where $\pi_{1}$ is a projection onto the first coordinate. This projection restricted to the graph of $g$ is a homeomorphism, so from the fact that $M\cap Gr(g)$ is analytic (co-analytic) we obtain that $B=\pi_{1}(M\cap Gr(g))$ is also analytic (co-analytic), hence it has the property of Baire. A contradiction with Lemma 7.16. ∎
8. Endpoints and absolute endpoints of chainable continua
Let $X$ be a chainable continuum. Recall that a point $x\in X$ is an endpoint of $X$ if for every $\varepsilon>0$, there is an $\varepsilon$-chain covering $X$ such that only the first link of the chain contains $x$. The following fact was proved in [Bi2].
Lemma 8.1 (Section 5 in [Bi2]).
Let $X$ be a chainable continuum. Then a point $x\in X$ is an endpoint of $X\iff$ the condition “if each of two subcontinua of $X$ contains $x$, then one of the subcontinua contains the other" is satisfied.
Following Rosenholtz [Ro], we say that a point $x\in X$ is an absolute endpoint of $X$ if for every $\varepsilon>0$ there is $\delta>0$ such that if $\{d_{1},...,d_{n}\}$ is a $\delta$-chain covering $X$ and $x\in d_{k}$, then $\bigcup_{j=1}^{k}d_{j}\subseteq B(x,\varepsilon)$ or $\bigcup_{j=k}^{n}d_{j}\subseteq B(x,\varepsilon)$, where $B(x,\varepsilon)$ is an open ball around $x$ of radius $\varepsilon$.
Below we indicate the relationships between the endpoints and the absolute endpoints of chainable continua and the minimal or maximal points of chainable continua equipped with ultrafilter orders.
Theorem 8.2.
Let $X$ be a chainable continuum and let $x\in X$. Then
  1. (1)
    $x$ is an endpoint of $X\iff$ there exists an ultrafilter order $\leq_{\mathcal{U}}^{\mathcal{D}}$ on $X$ such that $x$ is minimal or maximal with respect to $\leq_{\mathcal{U}}^{\mathcal{D}}$,
  2. (2)
    $x$ is an absolute endpoint of $X\Longrightarrow$ for every ultrafilter order $\leq_{\mathcal{U}}^{\mathcal{D}}$ on $X$ point $x$ is minimal or maximal with respect to $\leq_{\mathcal{U}}^{\mathcal{D}}$.
Proof of $(1)$.
$\Longrightarrow$ Suppose that $x$ is an endpoint of a chainable continuum $X$. Then for $n\in\mathbb{N}$ we can take a chain $D_{n}$, covering $X$, such that $x\in d_{1,n}$ and $mesh(D_{n})<\frac{1}{n}$. Let $\mathcal{D}=\{D_{n}\}_{n\in\mathbb{N}}$ be a sequence of such chains and let $\mathcal{U}$ be a non-principal ultrafilter on $\mathbb{N}$. Then we can observe that $x$ is a minimal point in the ultrafilter order $\leq_{\mathcal{U}}^{\mathcal{D}}$.
[…]
Proof of (2): Suppose that $x$ is an absolute endpoint in $X$ and let $\leq_{\mathcal{U}}^{\mathcal{D}}$ be an ultrafilter order on $X$. By [Ro, Theorem 1.0 (2)], $X\setminus\{x\}$ is a composant of X. Since composants are intervals in the linearly ordered space $(X,\tau_{\mathcal{U}}^{\mathcal{D}})$ (this follows from Lemma 4.2), we know that $X\setminus\{x\}$ is a subinterval of $(X,\tau_{\mathcal{U}}^{\mathcal{D}})$, hence $x$ must be maximal or minimal in $(X,\leq_{\mathcal{U}}^{\mathcal{D}})$. ∎
We do not know if the implication in the part (2) of the above theorem can be reversed, see Question 9.6.
Remark 8.3.
[…]
For $1\leq i<j$, we denote the bonding map $f_{i}\circ f_{i+1}\circ\cdots\circ f_{j-1}:I_{j}\to I_{i}$ by $f^{j}_{i}$, and $p_{i}:X\to I_{i}$ denotes the projection onto $i$th axis.
Note that the point $(0,0)$ in $K$ corresponds to the point $(0,0,0,\dots)$ in $X$.
Fix $i\geq 1$ and let $x^{i}_{k}=k2^{-(2i+1)}$ for $k=0,1,\dots 2^{2i+1}$.
We take $r=(2/3)|x^{i}_{k}-x^{i}_{k+1}|=(1/3)2^{-2i}$ and define the following sequence of open subintervals of $[0,1]$:
$O^{i}_{0}=[0,r),O^{i}_{1}=(x^{i}_{1}-r,x^{i}_{1}+r),O^{i}_{2}=(x^{i}_{2}-r,x^{i}_{2}+r),\dots O^{i}_{2^{2i+1}}=(1-r,1]\,.$
Put
$V^{i}_{0}$ $=$ $O^{i}_{2^{2i-1}};$
$V^{i}_{k}$ $=$ $O^{i}_{2^{2i-1}-k}\cup O^{i}_{2^{2i-1}+k}\quad\mbox{for }k=1,2,\dots 2^{2i-1}-1;$
$V^{i}_{2^{2i-1}}$ $=$ $O^{i}_{0}\cup O^{i}_{2^{2i}}\cup O^{i}_{2^{2i+1}};$
$V^{i}_{2^{2i-1}+k}$ $=$ $O^{i}_{2^{2i}+k}\cup O^{i}_{2^{2i+1}-k}\quad\mbox{for }k=1,2,\dots 2^{2i-1}-1;$
$V^{i}_{2^{2i}}$ $=$ $O^{i}_{2^{2i}+2^{2i-1}}.$
Let $U^{i}_{k}=p^{-1}_{i+2}(V^{i}_{k})$, for $k=0,1,\dots 2^{2i}$ and $E_{i}=(U^{i}_{0},U^{i}_{1},\dots U^{i}_{2^{2i}})$.
Let $d$ be the metric in $X$ defined by the formula
$d((x_{i}),(y_{i}))=\max_{i}2^{-i}|x_{i}-y_{i}|\quad\mbox{for }(x_{i}),(y_{i})\in X.$
Observe that, for $V^{i}_{k}$ of the form $O^{i}_{l}\cup O^{i}_{m}$ or $O^{i}_{l}\cup O^{i}_{m}\cup O^{i}_{n}$, we have $f^{i+2}_{i}(V^{i}_{k})=f^{i+2}_{i}(O^{i}_{l})$. Since $\operatorname{diam}(O^{i}_{l})\leq 2r$, one can easily calculate that, for $j=1,2,\dots i$,
$\operatorname{diam}(f^{i+2}_{j}(V^{i}_{k}))=\operatorname{diam}(f^{i+2}_{j}(O^{i}_{l})\leq 2r2^{i+2-j}\leq r2^{i+1}=(1/3)2^{-i+1}\,.$
This estimate implies that $mesh(E_{i})$ tends to 0.
A routine verification shows that for $\mathcal{E}=(E_{i})_{i}$ and any ultrafilter $\mathcal{V}$ we have the following ordering
$(1,1/2,1/4,1/8,\dots)\leq_{\mathcal{V}}^{\mathcal{E}}(0,0,0,0,\dots)\leq_{\mathcal{V}}^{\mathcal{E}}(1,1/2,3/4,5/8,\dots),$
where $(1,1/2,3/4,5/8,\dots)$ is a uniquely determined point $(x_{i})$ of $X$ such that $x_{1}=1$ and $x_{i}\geq 1/2$ for all $i$. This shows that the point $(0,0,0,\dots)$ is neither minimal nor maximal with respect to the order $\leq_{\mathcal{V}}^{\mathcal{E}}$.
Let $C$ be the composant of $X$ containing the point $(0,0,0,\dots)$. One can easily check that $(0,0,0,\dots)$ is a maximal point in $C$ with respect to the order $\leq_{\mathcal{V}}^{\mathcal{E}}$.
Since the composants of the Knaster continuum coincide with the arc components, by Corollary 4.4 we know that for every composants $C_{1},C_{2}$ of $X$ we have
[…]
Note that there are two cases:
  1. (1)
    for every composant $C^{\prime}$ of $X$ such that $C\leq_{\mathcal{V}}^{\mathcal{E}}C^{\prime}$ there exists a composant $D$ in $X$ such that C ≤_V^E D ≤_V^E C’,
  2. (2)
    there is a composant $C^{\prime}$ in $K$ such that $C\leq_{\mathcal{V}}^{\mathcal{E}}C^{\prime}$ and there is no composant $D$ in $K$ such that C ≤_V^E D ≤_V^E C’  . [^1] Actually, we can prove that the case (2) is not possible, but our proof is much longer than the above discussion of this case.
We observe that in both above cases the space $(K,\tau_{\mathcal{U}}^{\mathcal{D}})$, where $\tau_{\mathcal{U}}^{\mathcal{D}}$, is the order topology generated by an ultrafilter order on $K$ described in Subsection 6.1, is not homeomorphic to the space $(X,\tau_{\mathcal{V}}^{\mathcal{E}})$.
  • In the case $(1)$ the space $(X,\tau_{\mathcal{V}}^{\mathcal{E}})$ is not locally connected at the point $(0,0)$ and the space $(K,\tau_{\mathcal{U}}^{\mathcal{D}})$ is locally connected.
  • In the case $(2)$ we have:
    $(X,\tau_{\mathcal{V}}^{\mathcal{E}})\stackrel{{\scriptstyle\text{homeo}}}{{\simeq}}\bigoplus_{j\in J}Y_{j},$
    where all $Y_{j}$ are homeomorphic to the open interval $(0,1)$. By Theorem 6.4 spaces $(X,\tau_{\mathcal{V}}^{\mathcal{E}})$ and $(K,\tau_{\mathcal{U}}^{\mathcal{D}})$ are not homeomorphic.
Since $X$ is a topological copy of the Knaster continuum $K$, we conclude that $K$ admits two ultrafilter orders such that the order topologies on $K$ generated by these orders give us nonhomeomorphic spaces.
@@ 9. Questions
$x\leq_{\mathcal{U}}^{\mathcal{D}}y\iff h^{\prime}(x)\leq_{\mathcal{V}}^{(I_{i},f_{i})_{i=1}^{\infty}}h^{\prime}(y)$
@@
One can easily show that ultrafilter orders on the inverse limit of arcs (in the sense of Definition 3.5) are dense. We conjecture that ultrafilter orders defined using sequence of chains are also dense orders.
Question 9.2.
@@
Again, one can show that the above question has an affirmative answer for ultrafilter orders in the sense of Definition 3.5.
Question 9.4.
@@
By Corollary 6.5 and Remark 8.3 we know that order topologies generated by certain ultrafilter orders on the Knaster continuum are non-connected and non-compact. We would like to ask if any order topology on any indecomposable chainable continuum also has those properties.
Question 9.5.
@@
Question 9.6.
Let $X$ be a chainable continuum and let $x$ be a point of $X$ which is not an absolute endpoint. Does there exist an ultrafilter order on $X$ such that $x$ is neither minimal nor maximal with respect to this order?
Problem 9.7.